A side and a longer diagonal of rhombus measure 100 m and 160 m respectively.
a. Find the length of the shorter diagonal.
b. Find the area of the rhombus.

Answer by mananth(16946)

ODC is a right triangle
CD is the hypotenuse
100^2 = 80^2+OC^2
OC^2= 100^2-80^2
OC^2 = 3600
OC =60
shorter diagonal = 2*60 =120
Area of rhombus = 1/2 * product of diagonals
= 1/2 * 120 *160
=9600 m^2